In a study of speed​ dating, male subjects were asked to rate the attractiveness of their female​ dates, and a sample of the results is listed below ​(1equalsnot ​attractive; 10equalsextremely ​attractive). Construct a confidence interval using a 95​% confidence level. What do the results tell about the mean attractiveness ratings of the population of all adult​ females? 8​, 7​, 3​, 9​, 6​, 5​, 7​, 8​, 7​, 10​, 5​, 8 What is the confidence interval for the population mean mu​?

What does the confidence interval tell about the population of all adult​ females? Select the correct choice below​ and, if​ necessary, fill in the answer​ box(es) to complete your choice.

A.We are 95% confident that the interval from nothing to nothing contains the true mean attractiveness rating of all adult females.(Round to one decimal place as​ needed.)

B.The results tell nothing about the population of all adult​ females, because participants in speed dating are not a representative sample of the population of all adult females.

C.We are confident that 95​%of all adult females have attractiveness ratings between nothing and nothing.

​(Round to one decimal place as​ needed.)
Answer #1

sample mean, xbar = 6.92
sample standard deviation, s = 1.9287
sample size, n = 12
degrees of freedom, df = n - 1 = 11

Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 2.2

CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (6.92 - 2.2 * 1.9287/sqrt(12) , 6.92 + 2.2 * 1.9287/sqrt(12))
CI = (5.7 , 8.14)

C.We are confident that 95​%of all adult females have attractiveness ratings between 5.70 and 8.14